The radius of the Sun is about 7 x108m. If a photon could travel from the core to the surface of the Sun without being hindered by the convection zone, how long would it take? (2 pts) You detect a star and determine its mass to be four times that of the Suns. What is your prediction for the ultimate fate of this star and why? Your options are Red Giant, White Dwarf, Neutron Star, and Black Hole. (2 pts)
Answer #1

SOLUTION

Photon travels with the speed of light.

Here, distance, d = 7*108 m

speed of light, c = 3*108 m/s

Time taken to travel to the surface, t = d/c = 7/3 = 2.3 sec

If the mass of the star is four times that of the sun then after collapsing it becomes neutron star.

A star is basically just hot gas, the only thing that is supporting it under gravity is its temperature. So, it slowly cools down and shrinks and forms this very big lump of helium as sort of helium White Dwarf. Normal mass stars, like the Sun, burn their hydrogen away but then they've got enough mass to collapse down further and begin burning their helium to form carbon; then that will burn away. As it does so, the core of the star collapses. It gets very, very hot and blows the outer layers of the star out to form a red giant. It's got a very small core with a great big kind of diffuse, warmish red star outside it. This core is not massive enough to burn the carbon to form anything else. So, the core can make an explosion, it blows away the gas to leave a cold, carbon core.

Some of these, as they cool down, can crystallize and form diamond-type things. If you get a bit bigger than this white dwarf, that's got enough mass to collapse and form a neutron star, that explodes and forms a huge supernova. If you get even bigger than that, it's so massive that the neutron star will collapse to form a black hole from which nothing even light can escape.